### v2

# Inferring – Evolution from the Gravitational Wave Background

Abstract
We test the impact of an evolving supermassive black hole (SMBH) mass scaling relation (–) on the predictions for the gravitational wave background (GWB). The observed GWB amplitude is 2-3 times higher than predicted by astrophysically informed models which suggests the need to revise the assumptions in those models. We compare a semi-analytic model’s ability to reproduce the observed GWB spectrum with a static versus evolving-amplitude – relation. We additionally consider the influence of the choice of galaxy stellar mass function on the modeled GWB spectra. Our models are able to reproduce the GWB amplitude with either a large number density of massive galaxies or a positively evolving – amplitude (i.e., the ratio was higher in the past). If we assume that the – amplitude does not evolve, our models require a galaxy stellar mass function that implies an undetected population of massive galaxies ( at ). When the – amplitude is allowed to evolve, we can model the GWB spectrum with all fiducial values and an – amplitude that evolves as .
show]cayenne@umich.edu
1 Introduction
In mid-2023, four pulsar timing array (PTA) collaborations announced evidence for the gravitational wave background (GWB) (R. W. Hellings & G. S. Downs, 1983; G. Agazie et al., 2023a; EPTA Collaboration et al., 2023; D. J. Reardon et al., 2023; H. Xu et al., 2023). In all four announcements, the strain amplitude of the signal was 2–3 times greater than expected. It is commonly thought that the source for this quadrupolar signal is dominated by supermassive black hole (SMBH) binaries (M. C. Begelman et al., 1980; M. Milosavljević & D. Merritt, 2001; S. Burke-Spolaor et al., 2019). In this case, the GWB amplitude is most sensitive to the chirp mass of the binaries and is therefore expected to be dominated by the most massive binaries (E. S. Phinney, 2001). Thus, to model the GWB, it is necessary to have a reliable way to characterize the mass distribution of the underlying SMBH binary population. To make meaningful astrophysical inferences from PTA data it is therefore important to have an accurate predictor for the black hole mass function (BHMF), especially for the most massive SMBHs.
In the local universe, it has been shown that there is a strong correlation between host galaxy bulge mass and central SMBH mass, making it possible to accurately predict SMBH masses in nearby galaxies (J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). Direct, independent SMBH mass measurements are difficult outside the local universe, however, and so very few constraints exist for this relation for . High redshift SMBH mass measurements are hindered by telescope resolution limits meaning that masses must be inferred indirectly. Because of their extreme luminosity, AGN can be used to place useful constraints on SMBH masses at redshifts as high as (J. Jeon et al., 2025; Á. Bogdán et al., 2024; R. Maiolino et al., 2024a; L. Napolitano et al., 2024). From the AGN luminosity function, one can infer the BHMF, but assumptions about, e.g., radiative efficiency and AGN fraction lead towards large uncertainties (X. Shen et al., 2020). Furthermore, these types of surveys are frequently magnitude limited and are thus subject to observational biases (T. R. Lauer et al., 2007).
While AGN studies inform SMBH populations, modeling the GWB calls for a simpler BHMF framework. The observed correlation between galaxy bulge mass and SMBH mass is known as the – relation (J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). To approximate SMBH number density at higher redshifts, one can convolve the local – relation with the galaxy stellar mass function (GSMF), which is well-observed for the redshifts where most of the GWB signal is expected to be produced ( A. Sesana et al., 2004; G. Agazie et al., 2023b; J. Leja et al., 2020). This method assumes the – relation to be unchanging with time, an assumption with ambiguous observational support. It expected that analyses of the GWB can be used to place constraints on the underlying population of galaxies and black holes (e.g., J. Simon & S. Burke-Spolaor, 2016; D. Izquierdo-Villalba et al., 2022; S. Bonoli et al., 2025). G. Agazie et al. (2023b) derived physical quantities from fitting the NANOGrav 15-year PTA data and their results suggested that the data are best described with a number density of SMBHs, with , that is notably greater than predictions from current models (such as J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). Since then, other studies have used a variety of techniques finding similar results (e.g., G. Sato-Polito et al., 2024; Y. Chen et al., 2024; E. R. Liepold & C.-P. Ma, 2024). Additionally, there exist other scaling relations, such as the – relation, which connects SMBH mass to galaxy velocity dispersion (L. Ferrarese & D. Merritt, 2000; K. Gebhardt et al., 2000). It has been shown that the – and – relations make different predictions for the BHMF at high redshift such that the predicted GWB is higher when using – (C. Matt et al., 2023; J. Simon, 2023). This difference in BHMF predictions may be due to evolution in one or both relations that is not accounted for. Via the three-way relationship between galaxy stellar mass, radius, and velocity dispersion (A. de Graaff et al., 2021), the – relation naturally accounts for galaxy downsizing (A. van der Wel et al., 2014) and may therefore be a more fundamental probe of the SMBH’s gravitational potential well than – (R. C. E. van den Bosch, 2016; J. H. Cohn et al., 2025). This combination of findings suggests that methods of predicting SMBH number density are in need of revision.
It is unknown when the – relation was first established or how accurate it is outside the local universe. The strong local correlation between SMBH mass and galaxy bulge mass suggests the growth of these objects is coupled (J. Kormendy & L. C. Ho, 2013). If the – relation were to be unchanging with time, it would suggest a lockstep growth for galaxies and SMBHs. In other words, galaxy star formation matches the pace of SMBH accretion and / or the dominant growth mechanisms for both of these bodies are through major mergers or positive AGN/stellar feedback. It is unknown, however, whether SMBHs grow faster than their host galaxies, vice versa, or if they grow symbiotically over cosmic time. Whether a galaxy’s SMBH is over- or undermassive relative to the – relation throughout time has implications for the interactions between feedback from AGN accretion and galactic star formation (M.-Y. Zhuang & L. C. Ho, 2023; J. H. Cohn et al., 2025). Constraints on these growth pathways are necessary for determining SMBH number density and galaxy evolution outside the local universe.
Results from the James Webb Space Telescope (JWST) have found galaxies at high redshift () with overmassive black holes entirely unpredicted by the local – relation (e.g, Y. Harikane et al., 2023; F. Pacucci et al., 2023; J. Matthee et al., 2024; I. Juodžbalis et al., 2025). These high-redshift black holes are unexpectedly high in both mass and number density, with candidates up to (e.g. R. Maiolino et al., 2024b). Furthermore, there is an abundance of observational and simulation-based studies that have found evidence both for and against an evolution in the amplitude of the – relation (e.g., J. S. B. Wyithe & A. Loeb, 2003; A. Cattaneo et al., 2005; C. Y. Peng et al., 2006; D. J. Croton, 2006; P. F. Hopkins et al., 2009; K. Jahnke et al., 2009; R. Decarli et al., 2010; A. Merloni et al., 2010; B. Trakhtenbrot & H. Netzer, 2010; V. N. Bennert et al., 2011; M. Cisternas et al., 2011; X. Ding et al., 2020; J. Li et al., 2021; Y. Zhang et al., 2023; F. Pacucci et al., 2023; B. A. Terrazas et al., 2024; M. Yue et al., 2024; Y. Chen et al., 2024; A. P. Cloonan et al., 2024; M. R. Sah & S. Mukherjee, 2024; M. R. Sah et al., 2024; T. Shimizu et al., 2024; F. Pacucci & A. Loeb, 2024; A. Hoshi et al., 2024; M. M. Kozhikkal et al., 2024; Y. Sun et al., 2025; T. S. Tanaka et al., 2025).
An – amplitude that changes with cosmic time would have a significant impact on GWB predictions and interpretations. For example, if SMBH growth generally outpaces galaxy growth at higher redshifts, the – relation would therefore underestimate SMBH masses which would, in turn, lead to an underestimate of the GWB amplitude. Past studies have used electromagnetic (EM) observation, theory, and simulations to constrain potential evolution in the – relation, and now gravitational waves offer a new, independent, basis for testing.
If, instead, the – relation that is measured locally has not changed significantly since , a higher number density of massive galaxies in this redshift range could explain the high GWB amplitude. The galaxy stellar mass function (GSMF) is not observed to undergo significant evolution from to (e.g., J. Leja et al., 2020) though theory predicts otherwise (e.g., S. Tacchella et al., 2019). Recently E. R. Liepold & C.-P. Ma (2024) found that this lack of observed evolution may be an artifact of the survey design. They explain that, locally, the most massive galaxies are rare enough that most current integral-field spectroscopic surveys do not observe a large enough volume to catch them. This leads to an incomplete local sample for galaxies with masses . Using the MASSIVE survey (which has a field of view, C.-P. Ma et al., 2014) E. R. Liepold & C.-P. Ma (2024) measured a new GSMF based on this sample and found that the number density of these massive galaxies is significantly higher than other measurements (such as M. Bernardi et al., 2013; J. Moustakas et al., 2013; R. D’Souza et al., 2015; J. Leja et al., 2020). Their work suggests that predictions for the GWB may have previously been under counting the number of the most massive galaxies, and therefore SMBHs, which could lead to an underestimate of the GWB amplitude. This intriguing possibility warrants further testing.
In this paper we evaluate the possibility of an evolving – amplitude and test the impact of this evolution on the GWB spectrum. We additionally consider changes to the GSMF and explore the degeneracy between these two solutions. In section 2 we describe our model setup including our functional form of redshift evolution of the – relation. In section 3 we present the results of our models. We discuss the implications of our work in section 4, and a summary of our work and conclusions can be found in section 5.
2 Methods
In this section we describe the general setup of our semi-analytic model and our choices for the – scaling relation and GSMF. We additionally provide details for the priors and assumptions for our different test cases in section 2.4.
2.1 Semi-Analytic Modeling
We use holodeck (details can be found in section 3 of G. Agazie et al., 2023b)111Also see https://github.com/nanograv/holodeck. to synthesize populations of massive black holes. We start by convolving the galaxy stellar mass function from J. Leja et al. (2020) with the SMBH–galaxy mass scaling relation from (J. Kormendy & L. C. Ho, 2013). The specifics of this implementation are detailed in sections 2.2 and 2.3. We then apply galaxy merger rates, and SMBH binary hardening models to solve for the number density of SMBH binaries emitting gravitational waves in frequencies detectable by PTAs, i.e., the GWB spectrum predicted from the input parameters. Previously, G. Agazie et al. (2023b) calculated galaxy merger rates from galaxy pair fractions and merger times following the prescription due to (S. Chen et al., 2019a). For this work, we instead use the galaxy merger rate prescription due to V. Rodriguez-Gomez et al. (2015), which provides a more astrophysically motivated model for mergers. This change does not significantly affect the shape or amplitude of the GWB spectrum (see Figure 10 in Appendix A). A detailed description and graphic of the holodeck workflow can be found in section 3 of G. Agazie et al. (2023b).
To determine our best-fit parameters for each model, we first sampled 20,000 times from a set of astrophysically-motivated prior distributions with a Latin hypercube. We generated and averaged over 100 realizations for each set of sampled parameters to account for the Poisson sampling in the GWB spectrum calculations. Our sample size is higher than was used in G. Agazie et al. (2023b), who used Gaussian process interpolation for their posterior parameter estimation. Because of our larger parameter space (up to 30 parameters versus their 6), it is not computationally feasible to follow their same methods (N. Laal et al., 2025). Instead, our finer sampling makes it so that we can compare each model to the spectrum without the need for Gaussian process interpolation between the models. We compare our models to the 15 year Hellings and Downs (HD)-correlated free-spectrum representation of the GWB (W. G. Lamb et al., 2023). Specifically, we use the “HD-w/MP+DP+CURN” model which included both monopole (MP) and dipole (DP) correlated red noise in addition to the common uncorrelated red noise (CURN) (see additional details in G. Agazie et al., 2023b). For each GWB frequency, we evaluate the goodness of fit for a given modeled spectrum by comparing the value of each model to the probability density distribution of the data to determine how well each model fits the data. This process produces results with the same fidelity as G. Agazie et al. (2023b) with a higher computational efficiency (see their Appendix C, also our figures in Appendix C).
Throughout this paper we refer to the “likelihood” of the models which we calculate in our fitting process. These likelihoods represent how well models fit the GWB data, but do not contain information on how well the models agree with observational constraints from EM datasets. Therefore we additionally quantify goodness of fit to, e.g. the GSMF, with statistical tests such as the Kullback-Leilbler divergence (, S. Kullback & R. A. Leibler, 1951, see Section 3), and a quantity, , which we define in Section 3.4 to evaluate the agreement between our models and the observed GSMF.
2.2 The – relation
The local relation between SMBH and galaxy bulge mass can be described by a power-law relation (see, e.g., J. Kormendy & K. Gebhardt, 2001; J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). The value of the -intercept (amplitude) of this relation, throughout time, encodes the extent to which SMBHs and galaxies grow at the same rate. A constant amplitude value means that growth is tightly coupled and the coupling mechanisms are constant throughout cosmic time. In this work, we model a changing amplitude with a power-law evolution. We modified the existing – framework in holodeck to allow for this evolution by replacing the constant amplitude with one that is a function of redshift, . We parameterize it as
with
where the values can be determined from observation (; ; J. Kormendy & L. C. Ho, 2013), and can be positive or negative with indicating no evolution in the relation.
This power-law form of evolution is more rapid at lower redshifts, which means that the changes in the – relation are greatest in the redshift range most relevant to the GWB (, A. Sesana et al., 2004; G. Agazie et al., 2023b). We tested several functional forms of this evolution and determined that other options either evolve too slowly in the relevant redshift range, or are not distinguishable from a power-law with the current data. Future studies with higher signal-to-noise PTA data may eventually be able to place constraints on different functional forms, but until then, the power-law form is sufficient for this analysis. Throughout this analysis, we use the J. Kormendy & L. C. Ho (2013) values for the local – relation; we briefly discuss the impact of using alternative fits in Appendix A.
2.3 The Galaxy Stellar Mass Function
G. Agazie et al. (2023b) used the single-Schechter (P. Schechter, 1976) GSMF prescription due to S. Chen et al. (2019b). For this work, we predominantly use the GSMF due to J. Leja et al. (2020), which is defined as a double-Schechter function offering a more accurate description of the mass function of the total galaxy population (e.g., D. J. McLeod et al., 2021, and references therein). This GSMF is strongly supported by observational data and has an explicitly defined evolution making it possible to reliably calculate the GSMF at any specific redshift. This model is reliable within the range of the data () though we extrapolate the model to higher redshifts in holodeck for defining our populations. The reliability of extrapolating this model is uncertain, but the majority of the GWB signal is dominated by lower redshifts (G. Agazie et al., 2023b; A. Sesana et al., 2004) so we do not expect our results to be influenced by any error that may be introduced by this extrapolation. J. Leja et al. (2020) find that the redshift evolution of the characteristic mass and each of the density normalization terms is best described by a Gaussian. Following their methods (see their equations 14 and B 1-4), we define the full functional form of the GSMF to be
where
and
is the characteristic mass, is density normalization, and and are the upper and lower slopes of the power-law.
Recently, E. R. Liepold & C.-P. Ma (2024) developed a GSMF which has a notably higher number density than that of J. Leja et al. (2020) for . They make an estimate of the GWB amplitude and conclude that, using their local GSMF and an implied redshift evolution informed by the fractional GWB contribution functions from G. Agazie et al. (2023b) (see their Figure 12). Using their GSMF they claim they can produce a GWB amplitude that is consistent with what is seen by PTAs. Motivated by their results, we also consider how changes to the GSMF could instead explain the discrepancy between the predicted and observed GWB amplitude. E. R. Liepold & C.-P. Ma (2024) do not provide an explicit evolutionary form for their GSMF and so we assume an evolution that is equivalent to their GSMF and consistent with J. Leja et al. (2020) for . Further details of this evolution and discussion of alternative models can be found in Appendix A.
2.4 The Models
We present 8 models, which are summarized in Table 1. There are four different parameter setups and for each setup we ran one version of the model using the non-evolving – relation, and one version with evolution. For the models that allow for evolution of the – amplitude, we used a uniform prior for the evolutionary parameter .
The largest models include 29 parameters (30 with ). Of these parameters, the GSMFs make up 11 variables, five that describe the local GSMF and six that contain information about the GSMF evolution. For six of our eightmodels the priors for sampled GSMF parameters are based on the posterior fit values from J. Leja et al. (2020). Our fiducial parameters are therefore the median value of these distributions. For two of our eightwe use the same functional form of the GSMF, but with local values due to E. R. Liepold & C.-P. Ma (2024) and evolutionary parameters we define for this work (see appendix A). To help understand the degeneracy between a changing GSMF versus – relation we consider three different cases of sampling the J. Leja et al. (2020) GSMF: one in which we sample all 11 variables (Le11ne and Le11ev), one in which we only sample three local values (Le03ne and Le03ev), and one with all GSMF parameters fixed to their fiducial values (Le00ne Le00ev). The three GSMF values that are varied in Le03ne and Le03ev are , , and . For all but two of our 8 models we use the Gaussian posteriors from J. Leja et al. (2020) as priors for our GSMF and we assume that these parameters are independent of each other. For the last two models, we use our explicitly evolving version of the E. R. Liepold & C.-P. Ma (2024) GSMF, which we described previously in section 2.3 and also in appendix section A.
Each model we present is given a short-hand name to indicate the important features. Each name is six total characters to convey three pieces of information: The first two characters indicate which version of the GSMF is being used where Le is J. Leja et al. (2020) and LM is E. R. Liepold & C.-P. Ma (2024). The middle two characters indicate how many of the GSMF parameters (out of 11) were sampled where the number of parameters fixed to their fiducial values is 11 minus that number. The final two characters indicate whether the – amplitude was allowed to evolve (ev) or not (ne). This information is summarized in Table 1.
| Model Name | GSMF Parameters | Evolving – |
| Le11ne | 11 sampled | No |
| Le11ev | 11 sampled | Yes |
| Le03ne | 3 sampled | No |
| Le03ev | 3 sampled | Yes |
| Le00ne | 0 sampled | No |
| Le00ev | 0 sampled | Yes |
| LM11ne | 11 sampled | No |
| LM11ev | 11 sampled | Yes |
Because the models with evolving – relations are a strict parametric super set of those without (i.e., the models are nested), and since our prior for is uninformative, we performed Savage-Dickey ratio tests to quantify the significance of the evolving model over the fixed model (J. M. Dickey, 1971; E.-J. Wagenmakers et al., 2010). For all our figures and calculations we present the results from fits to the first 5 frequency bins. Our conclusions are not sensitive to the number of frequency bins we fit to.
3 Results
| Model Name | % Positive | S-D Ratio | |
| Le11ev | 0.84 0.79 | 87.2 | 1.3 |
| Le03ev | 1.02 0.48 | 99.3 | 24.6 |
| Le00ev | 1.05 0.55 | 96.8 | 5.2 |
| LM11ev | 0.86 0.76 | 88.7 | 1.5 |
A summary of the results for the posterior values for are shown in Table 2 and Figure 1. In Figure 2 we highlight parameters of interest for which we calculate Kullback-Leibler divergences, , between posterior and prior distributions (S. Kullback & R. A. Leibler, 1951). The best-fit spectra for all our models are presented in Figure 3. We find that, in general, a large number of massive SMBHs are needed to reproduce the GWB. When is allowed to vary, the majority of the posterior distribution is positive with a median value of across all models and for models with statistically significant evolution. We additionally find that, for models with fixed to be 0, this high number density of massive SMBHs can be modeled either through a top-heavy GSMF (between 0.2 to 3 dex higher in number density for and compared to the observed GSMF, see Figures 4 and 5) or a high – amplitude (with typically compared to the locally observed value of 8.69, J. Kormendy & L. C. Ho, 2013). We first present our full-sample model containing posterior distributions and best-fit spectra for all 29 (30 with ) parameters. Then, to highlight the effects of the degeneracy between parameters, we repeat this process with a subset of sampled parameters.
In section 3.4 we present our analysis of the GSMF parameters which we then further discuss in section 4.
3.1 Full-Sample Models
Here we present the results of models that sampled all 29 (30 with ) parameters. First we describe the outcome of the models that used our fiducial GSMF from J. Leja et al. (2020), then we present the impact of our test cases for explicitly evolving GSMF based on E. R. Liepold & C.-P. Ma (2024).
3.1.1 Le11ne and Le11ev
For the majority of parameters, the posterior distributions recover the priors for both Le11ne and Le11ev. In the top row of Figure 2 we highlight eight parameters from these models; we discuss the effects of the remaining parameters in appendix C and complete corner plots can be found in Appendix D. The non-evolving model, Le11ne shows slight deviations towards higher values for three of the the GSMF characteristic mass and normalization evolutionary parameters (, , and ). The two posterior distributions with the largest deviations from the priors are both of the evolutionary parameters for the characteristic stellar mass. When the – relation is allowed to evolve, the posterior distributions for Le11ev are more consistent with the priors for all GSMF parameters, though there is still a mild offset in the same direction as Le11ne for . The posterior distribution for is 87.2% positive with a median value of 0.85. A Savage-Dickey ratio test in favor of the evolving model returns a value of 1.3, which indicates that this fit is consistent with no – evolution.
The local GSMF values are recovered in both models, but the non-evolving model posteriors return generally larger values for the evolutionary parameters. Higher values for and produce greater number densities of galaxies at higher redshifts, while maintaining consistency with the fiducial local number density. In particular, larger values for the characteristic mass produce more massive galaxies and therefore a greater number of massive SMBHs.
We use to quantify the deviation of the posteriors from the priors for Le11ne and Le11ev. Two distributions with are equivalent while indicate disagreement. For all the parameters shown in Figure 2, the value of for Le11ne is equivalent or greater than for Le11ev. This means that the posterior distributions for Le11ne diverge from the priors more than in Le11ev suggesting that the posterior distributions for Le11ev are in better agreement with the observed GSMF and local – relation.
Figure 4 (see also Fig. 5) compares the GSMFs from J. Leja et al. (2020) to the GSMF implied by the median posterior values from Le11ne and Le11ev. In the local universe, all three GSMFs are in good agreement. By redshift 2, both GSMFs from Le11ne and Le11ev lie above the observed values, but Le11ev is consistent, within error bars with J. Leja et al. (2020). The difference between the GSMFs becomes larger with redshift with Le11ev remaining consistent with the observed GSMF while Le11ne is 0.5 to 2 orders of magnitude greater for galaxies with at . These massive galaxies are the hosts to the SMBHs in the mass range that dominates the GWB signal ( G. Agazie et al., 2023b). When predicting SMBH masses using the local – relation, to increase the amplitude of the modeled GWB, we need an increased number density of these massive galaxies compared to the fiducial model. Alternatively, an evolving – relation offers a way to increase the number density of the most massive SMBHs without any changes to the galaxy population or the locally observed SMBH–galaxy relation. This suggests that the best-fit parameters for models with an evolving – amplitude are more consistent with observational constraints than non-evolving models.
3.1.2 LM11ne and LM11ev
We find an anti-correlation between the strength of the GSMF evolution we assume and the resulting GWB amplitude. A strongly evolving GSMF (larger values of ) predicts lower mass densities of galaxies relative to a weakly evolving GSMF (lower values of ). The strongly evolving GSMFs also, therefore, produce lower number densities of the most massive SMBHs which then produces to lower GWB amplitudes.
Despite the greatly increased number density of local galaxies in this model, the strongly evolving GSMF produces a GWB spectrum that is only marginally greater in amplitude than the fiducial model. In fact, the posterior distributions, using the strongly evolving E. R. Liepold & C.-P. Ma (2024) GSMF (with its much higher number density of galaxies with ), display similar behavior to that of Le11ne and Le11ev. That is, the GSMF posteriors for models without – evolution tend towards greater number densities of high-mass galaxies for . The GSMF posteriors recover the priors when – evolution is allowed and the posterior distribution for is nearly identical to that of Le11ev (see Figure 1).
In Figure 5 we compare the prior and posterior GSMFs for all full-sample models. We see that the galaxy number densities for are roughly equivalent for Le11ne and LM11ne (similarly for Le11ev and LM11ev). We see a more significant difference between models that have an evolving or static – relation than between models with different prior GSMFs (further discussed in section 3.4). This result suggests that, to reproduce the GWB amplitude, we need a significantly greater number density of galaxies than we currently observe, not only locally, but at least out to . We discuss the physical implications and feasibility of this model further in section 4.
While the local GSMF measured by E. R. Liepold & C.-P. Ma (2024) offers valuable insight into the local galaxy population and the limits of volume-limited surveys, we do not find that this local increase in galaxy number density is sufficient to reproduce the observed GWB spectrum. The difference between our result and the what E. R. Liepold & C.-P. Ma (2024) find is a direct result of the assumptions made about binary hardening mechanisms in our respective calculations. For their estimate of the GWB amplitude, E. R. Liepold & C.-P. Ma (2024) follow the methods of G. Sato-Polito et al. (2024) and E. S. Phinney (2001). This model includes only the effect of gravitational wave-driven SMBH binary hardening, which produces a power-law spectrum. In this work, we assume that binaries can also harden through dynamical friction and stellar scattering. The effect of these additional hardening pathways is to change the shape of the overall spectrum and to lower the amplitude. While keeping all other parameters constant, the change in shape and amplitude has the greatest impact at the low-frequency end of the spectrum (, e.g., see lower-right panel of Figure 4 in G. Agazie et al., 2023a). The shape of the spectrum, when including these additional hardening mechanisms, provides better fits to the data across the observed range of frequencies than the gravitational wave-only model (G. Agazie et al., 2023a).
3.2 Fixed GSMF Evolution Models: Le03ne and Le03ev
The first of the sub-models that we discuss are Le03ne and Le03ev. Each of these is identical to Le11ne and Le11ev respectively except the six evolutionary parameters (, , , , , and ) alongside and for the GSMF are fixed to the posterior values given by J. Leja et al. (2020). To focus on the degeneracy between GSMF and – parameters we additionally fixed all parameters for galaxy merger rates and bulge fractions to their fiducial values. These models, therefore, sample a total of eight parameters (nine with ).
When sampling from all 11 GSMF parameters, the posteriors for the model that did not allow for – evolution (Le11ev) recovered the priors for the local GSMF values while only the evolutionary parameters deviated. Now, since these evolutionary parameters were fixed, the local GSMF posterior for characteristic mass is distributed towards greater values. Unlike before, this model also returns a posterior distribution with generally larger values of the – amplitude than the prior. An increase to either the characteristic galaxy mass or the – amplitude can generate the top-heavy BHMFs necessary to reproduce the observed GWB. A moderate (within 1) increase to both parameters, on the other hand, allows for massive SMBH populations to be modeled without deviating too far from either one parameter. This result is consistent with what G. Agazie et al. (2023b) found—to match the GWB either a large number of parameters in the fiducial model need to change by a small amount, or a very small number of parameters must be significantly different.
When the – relation is allowed to evolve, the posterior distributions for all GSMF parameters are in good agreement with the priors. Interestingly, the distribution for , for this model, is 99.3% positive with a Savage-Dickey ratio of 24.6. The standard deviation of this distribution is also significantly reduced indicating a higher confidence of positive evolution.
Similarly to Le11ne versus Le11ev, the between the priors and posteriors for Le03ne is equivalent or greater than for Le03ev. As mentioned earlier, the posterior distributions for and diverge from the priors more significantly in Le03ne than in Le11ne. In Le11ne the two distributions with the highest divergence were , and . These values are fixed in Le03ne and so this model compensates by increasing , and . We do not see a notable change to the values for these distributions in Le00ev versus Le11ev.
Nearly all samples from the model that did not include – evolution tend towards lower GWB amplitudes. In the second panel of Figure 3 we see that the shape of the GWB spectrum is nearly straight. Except for the two right-most frequency bins, the best-fit spectrum is equivalent to the upper limit on the models meaning that no models produced a GWB amplitude higher than the best-fit model for the higher-frequencies. The best-fit model here is roughly consistent (though on the low side) with four of the five frequency bins we used for fitting. The Le03ev model, on the other hand, while lower in GWB amplitude than Le11ev at the high frequency end, is still consistent with the PTA data.
3.3 Fixed GSMF Models: Le00ne and Le00ev
Finally, we present the smallest subset of our models, which only sample the three hardening parameters, the – amplitude, and (for Le00ev). These models fix all 11 GSMF parameters to their fiducial values from J. Leja et al. (2020) so the local and high- GSMF cannot vary for these models.
This subset of models returns a posterior distribution that is 96.8% positive with a Savage-Dickey ratio of 5.2. In this case, the posterior distribution for – amplitude recovers the prior almost exactly. For the non-evolving model, to reproduce high SMBH masses, a greater overall – amplitude is required. This result demonstrates how the models will trend more strongly towards positive values of as the other options for increasing the number density of massive SMBHs are removed. When the option for evolution is also removed, the posteriors will deviate strongly from the priors for parameters that are influential on the predicted SMBH number density.
The GWB spectra share similar characteristics to the Le03ne and Le03ev models with the non-– evolving models consistently lying below the majority of the PTA data and the evolving – models falling closer to the center of the violins in more bins, but lying slightly above the median. We note that the best-fit spectrum for Le00ne has the lowest likelihood of all 8 spectra indicating it is the worst fit to the GWB, an effect we detail further in Appendix C.
3.4 Influence of the GSMF
In this section we present the posterior GSMFs in two ways. We show the posterior GSMFs over-plotted with the corresponding prior GSMF for full GSMF sample models in Figure 5 for visual comparison. We then quantify the number density increase relative to J. Leja et al. (2020), which we plot versus the corresponding posterior median to demonstrate the degeneracy between these quantities in Figure 6.
In Figure 5 we can see that the posterior GSMF from Le11ne and LM11ne generally have higher number densities than the fiducial model. This offset is most extreme for the LM11ne GSMF which is to be expected since the fiducial GSMF for this model started with a higher number density than the J. Leja et al. (2020) GSMF for . Both of these GSMFs are greater in number density than their respective fiducial model and the fixed counterpart. Though we highlight the most extreme examples here, this is a trend across models that the fixed models require a greater number density of (especially the most massive) galaxies.
Because the GSMF is defined by up to 11 parameters in our models, the degeneracy between and any given GSMF parameter is not very strong. There is, however, considerable degeneracy between the – evolution and an increased number density of galaxies. To further demonstrate this degeneracy we define the quantity to be the integrated difference of the number density of galaxies between two models (see equation 6). We integrate over in mass and in redshift and note that trend of the results of this analysis are not sensitive to the bounds of our integral. We chose to omit in the integration range for clarity, including this range only results in a constant positive offset in for both LM11ne and LM11evẆe additionally normalize such that all our values lie between 0 and 1; indicates that the posterior GSMF is equivalent to the prior GSMF across our integration range (which is the case for both Le00ne and Le00ev). The parameter encapsulates the “boost” in galaxy number density and is analogous to the way encodes the changing – amplitude.
In Figure 6 we plot against the – amplitude at a fixed redshift , locally all values of are equivalent and the value of we choose only shifts the evolving – models left or right on the x-axis. The color differentiates evolving (red) models from non-evolving (blue) and the marker styles show the pairs of models with the same prior set up aside from – evolution (i.e., each marker style appears twice, one on red and one on blue). The clear negative trend demonstrates the degeneracy between an increased number density of galaxies and an increased SMBH–galaxy mass ratio in our posterior models. In the cases where the GSMF was allowed to vary, the evolving models (Le11ev, Le03ev, Le00ev, and LM11ev) have a lower value for and a higher value for than their non-evolving counterparts (Le11ne and Le03ne). These models follow the same trend, but are additionally boosted relative to the fiducial J. Leja et al. (2020) GSMF by design (as described in section 2.3 and Appendix A). The only outlier to this trend is Le00ne. The best-fit likelihood value for this model is several orders of magnitude smaller than for other models (represented by the size of the points) indicating it is a poor fit to the GWB data. While in good agreement with the GSMF (low ) its location on the plot is not an accurate reflection of the overall trend since the trend is defined by models that can reproduce the GWB. Therefore we conclude that the high number density of massive SMBHs, required to reproduce the GWB, can be produced either by a large increase to galaxy number density (GSMF), a large increase to the SMBH–galaxy mass ratio, or a moderate increase to both.
The models that fix some / all GSMF parameters converge to higher values of with a greater degree of confidence. The Le03ev model only fixed the evolutionary parameters from J. Leja et al. (2020) and sampled the local GSMF values. The posteriors for this model recovered the local GSMF and – parameters while its counterpart model, with fixed to 0, (Le03ne) did not. This suggests that the GWB free-spectrum is well described by a SMBH population as predicted from an evolving – relation and the galaxy number density given by the GSMF due to J. Leja et al. (2020). The GWB spectrum is equally well described by an unchanging – relation if, and only if, the GSMF is significantly higher in number density (for ), not only locally, but also out to as high as 3, which is not well supported by observational data (e.g., A. Muzzin et al., 2013; A. R. Tomczak et al., 2014; I. Davidzon et al., 2017; P. Behroozi et al., 2020), though it is not entirely ruled out (e.g., J. Moustakas et al., 2013; A. H. Wright et al., 2018).
3.5 Black Hole Number Density
In Figure 7 we present the BHMFs resulting from our models for . We additionally plot the posterior single SMBH functions from Figure 13 in G. Agazie et al. (2023b) and our fiducial BHMF for comparison. The BHMFs for evolving models using the J. Leja et al. (2020) GSMF (Le11ev, Le03ev, and Le00ev) are nearly identical to each other and so we plot one BHMF that is the median of all three models. We compare this to the posterior BHMF from LM11ev and find that the BHMFs are broadly consistent for , but that of LM11ev is greater in number density for the most massive SMBHs for . This is to be expected since this is the same behavior seen in the respective GSMFs.
We also show the median BHMF for the fixed models, Le11ne and Le03ne, (excluding Le00ne due to its low likelihood). This BHMF has a lower number density than that of the evolving models across the entire mass range for . This BHMF is more similar to the fiducial BHMF for low redshifts, but has a boosted number density which increases with redshift. This increasing positive offset is reflective of the GSMF posteriors and the increased number density of galaxies in the non-evolving best-fit models (for Le11ne and Le03ne).
Generally, all posterior models have a higher number density of SMBHs relative to the fiducial model across all redshift bins while evolving models typically have the highest number densities in a given bin.
4 Discussion
Our results suggest that either the – relation does not apply at higher redshift or galaxy observational surveys drastically underestimate the number of massive galaxies at all redshifts either through non-detection of a population or by underestimating masses. The best-fit evolving models indicate a positive offset for the – amplitude in the past. An – amplitude that increases with redshift () implies a black-hole-first evolution. Some studies have proposed this as a possible formation route stating that kinetic-mode feedback from SMBH accretion disks can impede the star formation within the host galaxy thus leading to a delayed stellar mass increase (M.-Y. Zhuang & L. C. Ho, 2023), also known as black hole “dominance” (M. Volonteri, 2012). In the opposite case, decreasing amplitude at high- may be indicative of radiative-mode feedback, where UV radiation from star formation prevents gas from sinking to the center of the galaxy therefore postponing SMBH accretion (M.-Y. Zhuang & L. C. Ho, 2023). Our results support the black hole dominance model of SMBH–galaxy coevolution. The value for the – amplitude evolution we find is consistent with several observational (X. Ding et al., 2020; F. Pacucci et al., 2023; Y. Zhang et al., 2023; T. S. Tanaka et al., 2025) studies and some theoretical / simulation studies (e.g., J. S. B. Wyithe & A. Loeb, 2003; A. Cattaneo et al., 2005; D. J. Croton, 2006), but there is a lack of consensus among these studies, especially for the redshifts most relevant for the GWB (e.g., A. Merloni et al., 2010; V. N. Bennert et al., 2011; J. Li et al., 2021; M. Cisternas et al., 2011). Our analysis places the strongest constraints only on the most massive SMBHs (). We cannot firmly differentiate between our power-law evolution and more complex models (such as dual sequence behavior T. Shimizu et al., 2024) nor can we determine whether the evolution is mass-dependent (e.g., A. Hoshi et al., 2024). A complementary follow-up analysis could consider the effect of heavy versus light SMBH seeding models on the low-mass end of the – relation to form a bigger picture of – evolution.
In Figure 8, we show the – amplitude offset () versus redshift predicted by our model alongside measurements from literature. The gray scattered points represent individual galaxies with both stellar and SMBH mass measurements and so we calculate the offset from the local – relation for these objects. The lines each represent a model from papers which provide an explicit functional form or offset for – evolution. The thicker red line and shaded red region represent the median posterior value for and 68% confidence region from our models with statistically significant evolution (Le03ev and Le00ev). We find that our model is most consistent with that from Y. Zhang et al. (2023) (yellow dotted line) though our error bars encompass the models due to J. S. B. Wyithe & A. Loeb (2003) and A. Merloni et al. (2010). We note that both M.-Y. Zhuang & L. C. Ho (2023) and A. Merloni et al. (2010) were careful to account for observational bias (e.g., as described by T. R. Lauer et al., 2007) in their analysis. Our model does not predict as extreme an offset as measured by X. Ding et al. (2020), F. Pacucci & A. Loeb (2024), or M. Yue et al. (2024).
We find which lies below the value () found by Y. Chen et al. (2024). Y. Chen et al. (2024) perform a similar analysis to the one we present here in which they fit the same data and also use the same fiducial – relation as us. There are two key differences in their model versus ours: (i) the GSMF and (ii) the SMBH binary hardening model. They use the GSMF due to P. Behroozi et al. (2020), which has notably different (both higher and lower) number densities to our fiducial model across the mass and redshift ranges we consider. Because the GSMF they assume is not uniformly offset from ours, the effect on their estimate is not obvious, but this difference is certainly a contributing factor to this discrepancy. Another difference between our models is the hardening prescription for the SMBH binaries. The model they use is a more complex model than we used here (Y. Chen et al., 2020). Differences in binary hardening efficiency can affect the GWB amplitude. We are able to model the GWB using if we assume a hardening timescale that is much longer than currently supported values (by up to a factor of ten times higher than reported in G. Agazie et al., 2023b). For this work, we assume a constant hardening timescale and the discrepancy between these results is demonstrative of the need for realistic binary hardening models. Future testing with EM studies will be important for placing bounds on hardening timescales. Inferences for SMBH populations from the GWB are sensitive to the underlying models used for galaxy populations. The difference in our results demonstrates the importance for having robust and well-constrained models.
Our analysis here suggests that, if we can reproduce the high GWB amplitude through only changes to the GSMF, then this would require a significantly higher number density of the highest mass galaxies at higher redshifts (at least out to ). As they discuss in their paper, the increased number density that E. R. Liepold & C.-P. Ma (2024) find indicates that galaxy surveys are missing out on the population of the most massive galaxies () and/or the mass to light ratios underestimate galaxy mass for this population. Local surveys may miss out on this population of galaxies because the survey area is simply not large enough to include this otherwise rare population. For higher redshifts, however, the survey volume is sufficient to ensure completeness and so the, e.g., COSMOS (C. Laigle et al., 2016) and 3D-HST+CANDELS survey (R. E. Skelton et al., 2014) surveys used in J. Leja et al. (2020) should be representative of the underlying population. This, combined with the abundance of reliable data, would suggest that the posterior GSMF at , needed by our models to match the GWB spectrum, is inconsistent with the true galaxy population. If, however, the initial mass function is bottom heavy, like that assumed by E. R. Liepold & C.-P. Ma (2024) and minimally evolving (and therefore bottom heavy at higher redshifts), then a GSMF like that in the bottom middle panel of Figure 5 could be feasible. In either case, analyses of the GWB are most directly a probe of SMBH properties, so any inferences we draw about the GSMF are several steps removed and are therefore only implicit. It is still useful to discuss the degeneracies with the GSMF in our models, however the GSMF estimates from this study should be treated with caution.
Gravitational wave-based studies, such as this, offer a new probe of SMBH–galaxy- mass scaling relations that can complement observational studies. Broadly speaking, there are two ways to observationally infer the SMBH mass function over cosmic time using only electromagnetic data: (i) BH mass scaling relations combined with observations of galaxies and (ii) models of Eddington ratio distributions combined with observations of AGN luminosity functions. These two methods have inferred incompatible SMBH mass functions, with galaxy-observation methods predicting higher densities than AGN-observation methods. A series of works (F. Shankar et al., 2016; E. Barausse et al., 2017; F. Shankar et al., 2019) proposed a potential solution whereby the (non-evolving) mass scaling relations had much lower amplitude (possibly a result of unsubstantiated measurement biases) to lower the galaxy-observation results to be in line with the AGN-observation results. Our work here shows that the GWB spectrum is generally in line with the galaxy-observation results. Not only are the measured scaling relations not too low at , they are, if anything, higher amplitude at higher redshift, something which is also noted by other studies (e.g., G. Sato-Polito et al., 2024; E. R. Liepold & C.-P. Ma, 2024). One as yet under-explored possibility to reconcile these two observational results is that radiative efficiencies are lower by a factor of 2 – 5 than typically assumed for the brightest AGN. This is outside the scope of this paper, but will be investigated in future studies.
Our analysis of the – relation focused on evolution of the amplitude, however it is possible that the relation may be evolving in other ways. Recent work has suggested that a variety of growth pathways are likely to be relevant within galaxy populations and that they are encoded in the intrinsic scatter (M.-Y. Zhuang & L. C. Ho, 2023; B. A. Terrazas et al., 2024; H. Hu et al., 2025; J. H. Cohn et al., 2025) where the local – relation acts as a sort of “attractor” for SMBH–galaxy pairs as they evolve. This concept is not new (e.g., C. Y. Peng, 2007; M. Hirschmann et al., 2010; A. Merloni et al., 2010; K. Jahnke & A. V. Macciò, 2011), but has seen a lot of new analysis, especially in the last two years. In particular, M.-Y. Zhuang & L. C. Ho (2023) conducted a study of AGNs and found that the scatter in the – relation is greater in the early universe, a finding substantiated recently by both gravitational wave and simulation analyses (e.g., E. C. Gardiner et al., 2023; B. A. Terrazas et al., 2024). Furthermore, J. Li et al. (2025) measured the SMBH–galaxy mass ratios for AGN at and found that they are consistent with the local population. Results such as theirs could be indicative of an increased scatter in the – relation outside the local universe. Biased observations of a high-scatter – relation could appear, artificially, as an inflated amplitude. Future GWB studies similar to this will provide valuable insights into alternative evolutionary forms of the – relation.
It is also possible that a scaling relation not based on galaxy mass may more accurately reproduce the highest-mass SMBH number density. It has been shown that the – and – relations predict different BHMFs outside the local universe (C. Matt et al., 2023) and that this has an impact on the predicted GWB amplitude (J. Simon, 2023). On the other hand, N. J. McConnell & C.-P. Ma (2013) found that both the – and – relations “saturate” towards the highest masses, underpredicting SMBH masses for the largest bright cluster galaxies. And an analysis by G. Sato-Polito et al. (2024) finds that neither of the local relations can reproduce the high GWB amplitude. In either case, there is strong evidence to suggest that the local – relation was different in the past in some way, and multimessenger studies such as this are an exiting route towards characterizing SMBH–galaxy coevolution.
5 Summary and Conclusions
In this work we implemented an evolving – model which improves our ability to reproduce the GWB while maintaining consistency with astrophysically constrained models. We adapted the holodeck semi-analytic model to test for evolution in the – relation and fit models to the GWB free-spectrum from G. Agazie et al. (2023a). Our results show mild to strong preference for a positive evolution in the – amplitude. When modeling the – relation as we find that .
We also studied the degeneracy between the GSMF and – relation. The GWB requires massive SMBHs and we find that this population can be modeled with a top-heavy GSMF (which becomes a top-heavy BHMF via the local – relation) or an increased – amplitude (either via positive redshift evolution or a high local value). The – relation remains the most influential component in GWB amplitude calculations. While an alternative GSMF offers an interesting solution, the GSMF has more robust observational constraints than the – relation outside the local universe. Moreover, the GWB is most directly a probe of SMBH properties and so any inference about the GSMF is a secondary calculation. We therefore conclude that a SMBH-first growth model provides the best fits to the GWB while maintaining a high degree of consistency with EM-based observational constraints.
Future work will investigate the effects of an evolving intrinsic scatter in the – relation. It will also be important to test different evolutionary models using upcoming PTA data releases which will have higher signal-to-noise to better differentiate between models.
Appendix A GSMF Evolution Parameters and Scaling Relation Choice
In this section, we describe the form of evolution we assume for the GSMF used in LM11ne and LM11ev. We additionally explore the effects of alternative GSMFs not presented in the main text and briefly investigate the implications for the GWB and high- galaxy population. The different GSMFs and corresponding GWB spectra are shown in Figure 9. We also describe the effect of – version on the predicted GWB spectrum at the end of this section and in Figure 10.
The GSMF measured by E. R. Liepold & C.-P. Ma (2024) uses only galaxies and so they do not provide an explicit evolutionary form for the GSMF at higher redshift. For use in our analysis, we construct and test several models that bracket the plausible rates of GSMF evolution. A strongly evolving GSMF is predicted for (e.g., G. De Lucia et al., 2007) though observations often show a lack of evolution in this redshift range (e.g., J. Moustakas et al., 2013; K. Bundy et al., 2017; J. Leja et al., 2020). The E. R. Liepold & C.-P. Ma (2024) GSMF offers observational support for this predicted evolution. The GSMF due to J. Leja et al. (2020) does not show significant evolution for which, according to E. R. Liepold & C.-P. Ma (2024) could be due to observational biases in the local universe. Large cosmological surveys do not have the appropriate local volume to catch the most massive galaxies and so the E. R. Liepold & C.-P. Ma (2024) GSMF offers to complete the picture for the local high-mass GSMF. The J. Leja et al. (2020) GSMF is consistent with many other studies and theory for , and there is no strong foundation to believe that the GSMF in this redshift range is invalid. We can assume that the J. Leja et al. (2020) GSMF is complete for and can therefore find an approximate evolution that matches the local E. R. Liepold & C.-P. Ma (2024) GSMF and also the J. Leja et al. (2020) for .
Starting with the explicit GSMF from E. R. Liepold & C.-P. Ma (2024) and assuming a similar functional form to that in J. Leja et al. (2020), we tested three different forms of evolution represented by the shaded regions in Figure 9: (i) Weakly evolving (green): producing a GWB amplitude similar to that in E. R. Liepold & C.-P. Ma (2024), (ii) Moderately evolving (orange), and (iii) Strongly evolving (blue): Designed to match the high-mass end of the J. Leja et al. (2020) GSMF by a redshift of , all parameters for these models are given in Table 3.
We believe the strongly evolving model to be the most faithful to the observational constraints from both E. R. Liepold & C.-P. Ma (2024) and J. Leja et al. (2020) hence this is the model we use for our analysis. The strongly evolving GSMF produces the lowest number densities at higher redshift compared to the other two evolutionary forms. The parameters for the strongly evolving GSMF were chosen such that the number density of massive galaxies is roughly similar to that of J. Leja et al. (2020) by redshift . This option bridges the gap between the local GSMF from E. R. Liepold & C.-P. Ma (2024) while maintaining consistency with J. Leja et al. (2020) and therefore represents the solution that is in agreement with both observational constraints. This is the model we use in LM11ne and LM11ev, the results of which are discussed in the main text in section 3.1.2.
The weakly evolving GSMF produces the highest GWB amplitude and is nearly consistent with the PTA data. In fact, only minor changes to, e.g., the hardening timescales would be needed to match the PTA data. This option represents the number density of galaxies that best reproduces the GWB starting from the local GSMF in E. R. Liepold & C.-P. Ma (2024) and without significantly changing any other assumptions in our model. We only provide this approximate form and do not perform any fits because this GSMF is unrealistic and only used for demonstration purposes.
The moderately evolving GSMF represents a sort of “compromise” between the other two extremes. Interestingly, the posterior GSMF from LM11ne is very similar to this toy model we present here. The change in number density across is actually greater than in J. Leja et al. (2020), though it is positively offset at every redshift compared to their GSMF. This GSMF could be a realistic representation of the high- population, but would suggest that we are vastly underestimating galaxy stellar masses at these redshifts (e.g., via an incorrect initial mass function) or that cosmological surveys are largely incomplete (by a factor of 10 or more) in this regime. At this time we cannot evaluate the feasibility of this GSMF though, as emphasized earlier in section 4, the inferences made for the GSMF from the GWB are to be treated with a high degree of caution.
| Evolution Type | |||||||||||
| Strong | |||||||||||
| Moderate | |||||||||||
| Weak |
Appendix B Choice of Priors and Prior Tests
We chose astrophysically motivated priors as inputs to all our models. We investigated the impact that prior choice has on the models and found that our results are not sensitive to prior choice.
Several of our parameters, e.g., galaxy merger timescales and SMBH binary hardening timescales, do not have strong observational constraints. Instead, we make use of theoretically motivated prescriptions. The two studies we base our priors on both make use of the Illustris cosmological hydrodynamical simulations (M. Vogelsberger et al., 2014). We use the galaxy merger rates due to V. Rodriguez-Gomez et al. (2015) which provides an analytical prescription for merger rates that depends on galaxy mass and redshift. This is a different model to that used in G. Agazie et al. (2023b) though we find that the impact on the resulting GWB spectrum is not significant. We ran a series of tests allowing only the galaxy merger rate parameters to vary and found that the posteriors always recovered the priors (and therefore the “best fit” model was identical to the prior model). We concluded that our models have minimal to no constraining power over these parameters hence they are fixed for Le03ne, Le03ev, Le00ne, and Le00ev.
Numerous studies have investigated evolution of the physical separation of SMBH binaries, i.e., the “hardening” rate (M. C. Begelman et al., 1980; G. D. Quinlan, 1996; G. D. Quinlan & L. Hernquist, 1997; Q. Yu, 2002; A. Sesana & F. M. Khan, 2015; L. Z. Kelley et al., 2017; A. K. Bhowmick et al., 2024; S. Buttigieg et al., 2025, Harris et al. in prep). SMBH binary hardening rates are difficult to constrain observationally and so these studies rely on hydrodynamic and/or N-body simulations. Modeling the entire SMBH merger process is computationally challenging, however, leaving large uncertainties in hardening models (L. Z. Kelley et al., 2017). For this work, we use uniform priors on all free hardening parameters in our models. We see that the posterior distributions often push up against the limits of our prior range. In our testing, we found that the qualitative results of this work were not sensitive to the prior ranges we assume for these parameters. For example the hardening timescale, lower limit is 0 Gyr and cannot go lower despite the posterior distribution pushing against this lower limit. Additionally the parameter lower limit is roughly at the limit at which the timescale of the stellar scattering regime is minimized and lower values of have a negligible impact on the GWB spectrum. Finally the third hardening parameter we sample is the characteristic radius, , which indicates the point at which the dominant hardening mechanism transitions from dynamical friction to stellar scattering. This radius corresponds to the sphere of influence of the SMBHs and is on the order of 10s of pc. Our priors range from 2–20 pc, which is physically motivated (L. Z. Kelley et al., 2017). Allowing for larger values of can be useful for studies focused on constraints at small binary separations, but these parameters are not the focus of this study so we chose a realistic, but uniform range of values. We also note that there is a degeneracy between and such that large values of correspond to larger values of . Therefore changing the prior range on may affect the posterior values for, e.g., , but it does not notably impact the results of GSMF or – parameters. Improving the hardening parameter prior choice for GWB studies requires a dedicated multimessenger study. This is outside the scope of this work, but will be included in future investigations (e.g., Harris et al. in prep).
For the GSMF and – relation, we used models based on electromagnetic observations (J. Leja et al., 2020; J. Kormendy & L. C. Ho, 2013; E. R. Liepold & C.-P. Ma, 2024). To evaluate the constraining power of our models over the GSMF, we ran a model sampling only the 11 GSMF parameters with uniform priors (Le11Une) and repeated this model with a free parameter (Le11Uev). We found that the results of this test were consistent with the results of our study. The posterior distributions deviated from the uniform priors and produced BHMFs that are consistent with each other and also the BHMFs from our models with informed priors. The BHMFs from these models are shown in Figure 11, we additionally over-plot the median BHMF from the Le_ne models from Figure 7. We see that all three BHMFs return remarkably similar number densities in all redshift bins though Le11Une falls slightly short of the two evolving – models, a pattern also present in Figure 7. The posterior – evolution parameter from Le11Uev is and the distribution is 82.8% positive. While this is not statistically significant evolution, we believe this result supports our overall conclusions and is a good indicator that our models are not prior dominated for the GSMF parameters.
Similar to how we tested different options for the GSMF, we additionally considered the effect of the version of the – relation we use. In Figure 10 we demonstrate the impact of scaling relation choice on the GWB spectrum. The version of the scaling relation from J. Kormendy & L. C. Ho (2013) generally produces bigger SMBHs (and therefore higher GWB amplitudes) than other options (see, e.g., J. Simon, 2023). E. R. Liepold & C.-P. Ma (2024) use the N. J. McConnell & C.-P. Ma (2013) version of the – relation, which generally produces lower GWB amplitudes compared to J. Kormendy & L. C. Ho (2013). We additionally show the GWB spectrum using the scaling relation from F. Shankar et al. (2016). This comparison demonstrates that, wherever our spectra fall short of the GWB data, using a relation with a lower amplitude would only exacerbate this problem. This also has implications for our choices for the evolution we approximate for the E. R. Liepold & C.-P. Ma (2024) GSMF.
Appendix C Reliability of Methods
Here we demonstrate that our “direct-likelihood” method reproduces the Gaussian process-interpolated MCMC-like methods of G. Agazie et al. (2023b). In Figure 12 we show a comparison between the posterior distributions from G. Agazie et al. (2023a) compared to our results when using the same prior model set up. We see that the two methods return negligibly different posterior distributions. The method we use in this work is also discussed in more depth in Appendix C of G. Agazie et al. (2023b).
In this work we discovered a minor limitation of the direct-likelihood method which impacted the Le00ne model. It is more difficult for the posterior distributions to deviate from the priors because extreme values (e.g. more than 3 above or below the prior median) have a low chance of being randomly sampled, regardless of how well that value may describe the data. We see this limitation in our model with the lowest number of sampled parameters, Le00ne. The best-fit GWB spectrum from this model returned a notably low log-likelihood compared to the other models which means it produces a worse overall fit (see Figures 3 and 6). We can see, from Le03ev and Le00ev, that a value of is needed to reproduce the GWB amplitude without increasing galaxy number density. If the – relation is not allowed to evolve, the prior for would need to be sampled at a value away from the prior mean (assuming a normal distribution where ), and therefore extremely unlikely / impossible to be sampled. Without interpolating, any value that is not sampled simply does not have a corresponding GWB spectrum and the best-fit parameters may therefore not describe the data. This effect can be mitigated by choice of priors, but prior choice has its own impact on models and so future studies should consider the impact of this effect when using this method.
It is important to emphasize that the likelihoods we calculate are a measure of how well a given best-fit spectrum describes the observed GWB spectrum, but they are not marginalized over the respective parameter spaces. This means that the likelihoods to not contain information about prior-posterior agreement/disagreement and therefore do not indicate goodness of fit to EM data. These likelihoods cannot be used alone to evaluate which models are better models of both the GWB and EM data. To evaluate consistency with EM data we make use of other quantities such as the Kullback-Leilbler divergence, , and GSMF boost, . The “best” models have high likelihoods (good fits to the GWB) in addition to low and values (good agreement with the EM data).
Appendix D Corner Plots
The full corner plots for each of our 8 models. Note that the variables pertaining to galaxy merger rates include “GMR” in the subscript to distinguish from other parameters which use similar variables. For the parameters in the model not detailed in this paper we refer the reader to G. Agazie et al. (2023b) (and references therein) for a complete discussion. In each corner plot, the blue histograms / contours represent the posterior distributions. The priors are shown by the gray histograms.
Fig. Set1. Model Output Corner Plots